# Recursive variational Gaussian approximation with the Whittle likelihood for linear non-Gaussian state space models

### v2

@@ 2.5 HMC-Whittle
Algorithm 4 One step of the Leapfrog algorithm, Leapfrog($\bm{\theta}^{*},\mathbf{r}^{*},\epsilon$)
@@
Set $\mathbf{r}^{*}=\mathbf{r}+\epsilon\nabla_{\bm{\theta}}\mathcal{L}\left({\bm{\theta}}\right)/2$,
Set $\bm{\theta}^{*}={\bm{\theta}}+\epsilon\bm{{M}}^{-1}\mathbf{r}^{*}$,
Set $\mathbf{r}^{*}=\mathbf{r}^{*}+\epsilon\nabla_{\bm{\theta}}\mathcal{L}\left(\bm{\theta}^{*}\right)/2$.
return $\bm{\theta}^{*}$, $\mathbf{r}^{*}$.
@@ 3 Applications
This section demonstrates the use of R-VGA-Whittle for estimating parameters of a univariate linear Gaussian SSM and those of a univariate/bivariate stochastic volatility (SV) model using simulated and real data. We compare the posterior densities obtained from R-VGA-Whittle to those from HMC-exact and HMC-Whittle. All examples are implemented in R, and the reproducible code is available from https://github.com/bao-anh-vu/R-VGA-Whittle. HMC-exact and HMC-Whittle are implemented using the CmdStanR interface to the Stan programming language (Gabry et al.,, 2024).
For all examples, we implement R-VGA-Whittle with damping and blocking as described in Algorithm 3. We apply damping to the first $n_{damp}=5$ angular frequencies with $D=100$ damping steps each. These values were chosen to reduce the initial instability at the expense of a small additional computational cost. We select $\tilde{n}$ for each model by finding the 3dB frequency cutoff and process the remaining frequencies in blocks of length $B=100$, following a study in Section S3 of the online supplement. At each iteration of R-VGA-Whittle, we use $S=1000$ Monte Carlo samples to approximate the expectations of the gradient and Hessian of the Whittle log-likelihood. This value of $S$ was chosen because it results in very little variation between repeated runs of R-VGA-Whittle; see Section S4 of the online supplement for an experiment exploring the algorithm’s sensitivity to $S$. HMC-Whittle and HMC-exact were run with two chains with 15000 samples each. The first 5000 samples of each chain were discarded as burn-in. All experiments were carried out on a high-end server, with an NVIDIA Tesla V100 32GB graphics processing unit (GPU) used to parallelise over the $S=1000$ Monte Carlo samples needed for estimating the expectations of the gradient and Hessian of the Whittle log-likelihood at each iteration of Algorithm 2.
@@ 3.1 Univariate linear Gaussian SSM
$T$ R-VGA-Whittle HMC-Whittle HMC-exact
Linear Gaussian 10000 14 125 447
SV (sim. data) 10000 13 104 17005
SV (real data) 2649 13 25 3449
2D SV (sim. data) 5000 121 14521 72170
2D SV (real data) 2649 107 8890 74462
Table 1: Computing time (in seconds) for R-VGA-Whittle, HMC-Whittle and HMC-exact for each model we considered. The number of observations $T$ for each example is also shown.
A comparison of the computing time between R-VGA-Whittle, HMC-Whittle and HMC-exact for this model can be found in Table 1. The table shows that R-VGA-Whittle is nearly 10 times faster than HMC-Whittle and more than 30 times faster than HMC-exact.
@@ 3.2 Univariate stochastic volatility model
SV models are a popular class of non-Gaussian financial time series models used for modelling the volatility of stock returns. An overview of SV models and traditional likelihood-based methods for parameter inference for use with these models can be found in Kim et al., (1998); an overview of MCMC-based inference for a univariate SV model can be found in Fearnhead, (2011).
In this section, we follow Fearnhead, (2011) and simulate $T=10000$ observations from the following univariate SV model:
@@
We set the true parameters to $\phi=0.99$, $\sigma_{\eta}=0.1$, and $\kappa=2$.
Following Ruiz, (1994), we apply a log-squared transformation,
@@
Since $\mathbb{E}(\log(y_{t}^{2}))=\mu$, the parameter $\mu$ can be estimated using the sample mean:
@@
$\hat{\kappa}=\sqrt{\exp(\hat{\mu}-\mathbb{E}(\log(\epsilon_{t}^{2})))},$
where $\mathbb{E}(\log(\epsilon_{t}^{2})))=\psi(1/2)+\log(2)$, and $\psi(\cdot)$ is the digamma function (Abramowitz and Stegun,, 1968). As we show below, the estimate $\hat{\kappa}$ is not required for R-VGA-Whittle and HMC-Whittle as the Whittle likelihood does not depend on $\kappa$. However, $\kappa$ appears in the likelihood used by HMC-exact. Since $\kappa$ is not estimated by the Whittle-based methods, for comparison purposes we fix it to its estimate $\hat{\kappa}$ when using HMC-exact.
@@
which leads to 95% probability intervals of $(0.57,0.99)$ for $\phi$, and $(0.11,0.44)$ for $\sigma_{\eta}$. The interval for $\phi$ aligns closely with those used in similar studies of the stochastic volatility (SV) model. For example, Kim et al., (1998) employed the transformation $\phi=2\phi^{*}-1$ with $\phi^{*}\sim\textrm{Beta}(20,1.5)$, resulting in a 95% probability interval for $\phi$ of $(0.59,0.99)$. Additionally, Kim et al., (1998) utilized an inverse gamma prior for $\sigma_{\eta}^{2}$ with shape and rate parameters of $2.5$ and $0.025$, respectively, corresponding to a 95% probability interval $\sigma_{\eta}$ for $(0.06,0.24)$, which is comparable to the interval implied by our chosen prior.
@@
Figure 5 shows the trajectories of the R-VGA-Whittle variational means for all parameters. The 3dB cutoff point for this example is the 75th frequency, after which there are 50 blocks, for a total of 125 updates. Similar to the case of the LGSS model in Section 3.1, the variational mean of $\phi$ is close to the true value from the early stages of the algorithm, while that of $\sigma_{\eta}$ requires more iterations to achieve convergence.
Figures 9(a)9(c) in Section S5 of the online supplement show the HMC-exact and HMC-Whittle trace plots. For this example, HMC-exact produces posterior samples with a much higher correlation than HMC-Whittle, particularly for the variance parameter $\sigma_{\eta}$. This is because HMC-Whittle only targets the posterior density of the parameters, while HMC-exact targets the joint posterior density of the states and parameters, which are high dimensional. We thin the posterior samples from HMC-exact by a factor of 100 for the remaining samples to be treated as uncorrelated draws from the posterior distribution.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle, and HMC-exact in Figure 6. The posterior densities from R-VGA-Whittle and HMC-Whittle are very similar, while the posterior densities from HMC-exact are slightly narrower. Table 1 shows that R-VGA-Whittle is 8 times faster than HMC-Whittle, and more than 1000 times faster than HMC-exact in this example.
@@
Figure 5: Trajectories of the variational means for parameters in the univariate SV model with simulated data. True parameter values are marked with dashed lines. The vertical line is at $\tilde{n}$, the point beyond which updates are done in blocks.
@@
Figure 6: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGAL-Whittle (red) and HMC-Whittle (yellow) for data simulated from the univariate SV model with $T=10000$. Diagonal panels: Marginal posterior distributions with true parameters denoted using dotted lines. Off-diagonal panels: Bivariate posterior distributions with true parameters denoted using the symbol $\times$.
@@ 3.3 Bivariate SV model
Figure 7: Trajectories of the variational means for parameters in the bivariate SV model with simulated data. True parameter values are marked with dashed lines. The vertical line is at $\tilde{n}$, the point beyond which updates are done in blocks.
@@
Figures 10(a)10(c) in Section S5 of the online supplement show the HMC-exact and HMC-Whittle trace plots for this model. HMC-exact samples are highly correlated, especially those of $\Sigma_{\eta_{22}}$. We find that it is necessary to thin the HMC-exact samples by a factor of 50 for the samples to be considered uncorrelated draws from the posterior distribution. In contrast, HMC-Whittle produces well-mixed posterior samples. Section S5 of the online supplement also shows that the effective sample sizes from HMC-Whittle are much higher than those from HMC-exact for all parameters.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle, and HMC-exact in Figure 8. In this example, we observe broad agreement between the three methods, though the posterior densities from R-VGA-Whittle and HMC-Whittle are wider than those from HMC-exact. The posterior density of $\Phi_{22}$ from R-VGA-Whittle puts more probability on values close to one than those from the other two methods, while the posterior density of $\Sigma_{\eta_{21}}$ from both R-VGA-Whittle and HMC-Whittle appear slightly shifted towards zero compared to that from HMC-exact. As shown in Table 1, R-VGA-Whittle is nearly 600 times faster than HMC-exact, and over 100 times faster than HMC-Whittle in this example.
@@
3.4 Real univariate and bivariate examples

### v3

@@ 2.5 HMC-Whittle
Algorithm 4 One step of the Leapfrog algorithm, Leapfrog($\bm{\theta},\mathbf{r},\epsilon$)
@@
Set $\mathbf{r}^{**}=\mathbf{r}+\epsilon\nabla_{\bm{\theta}}\mathcal{L}\left({\bm{\theta}}\right)/2$.
Set $\bm{\theta}^{*}={\bm{\theta}}+\epsilon\bm{{M}}^{-1}\mathbf{r}^{**}$.
Set $\mathbf{r}^{*}=\mathbf{r}^{**}+\epsilon\nabla_{\bm{\theta}}\mathcal{L}\left(\bm{\theta}^{*}\right)/2$.
Output: $\bm{\theta}^{*}$, $\mathbf{r}^{*}$.
@@ 3 Applications
This section demonstrates the use of R-VGA-Whittle to estimate parameters of various models: a univariate linear Gaussian state space model; univariate and bivariate stochastic volatility (SV) models; and a state space model with Student’s $t$ errors, where the latent states follow an autoregressive fractionally integrated moving average (ARFIMA) process. We evaluate R-VGA-Whittle on both simulated and real data, and compare the posterior densities obtained from R-VGA-Whittle to those from HMC-exact and HMC-Whittle. HMC-exact and HMC-Whittle are implemented using the CmdStanR interface to the Stan programming language (Gabry et al.,, 2024). For the univariate SV model in Sections 3.2 and 3.5, we also compare the posterior distributions from our algorithm to those obtained via the popular stochvol R package (Kastner and Frühwirth-Schnatter,, 2014). Reproducible R code for all examples is available from https://github.com/bao-anh-vu/R-VGA-Whittle.
For all examples, we implement R-VGA-Whittle with damping and blocking as described in Algorithm 3. Damping is applied to the first $n_{damp}=5$ angular frequencies with $D=100$ damping steps each. These values were chosen to reduce the initial instability at the expense of a small additional computational cost. We select $\tilde{n}$ for each model by finding the 3dB frequency cutoff and process the remaining frequencies in blocks of length $B=100$, following a study in Section S3 of the online supplement. At each iteration of R-VGA-Whittle, we use $S=1000$ Monte Carlo samples to approximate the expectations of the gradient and Hessian of the Whittle log-likelihood. This value of $S$ was chosen because it results in very little variation between repeated runs of R-VGA-Whittle; see Section S4 of the online supplement for an experiment exploring the algorithm’s sensitivity to $S$.
We run both the HMC-Whittle and HMC-exact algorithms using two parallel chains, each with a burn-in period of 1000 samples. The total number of iterations per chain is model-dependent, up to a maximum of 15000 samples (including burn-in). We summarise the number of posterior samples used in both HMC-exact and HMC-Whittle in Table S2 in Section S5 of the online supplement. These settings are chosen to target an effective sample size (ESS) of approximately $1000$ for all parameters. Detailed ESS results for each method are reported in Table S1 in Section S5 of the online supplement. Posterior distributions estimated using HMC-exact are treated as ground truth for comparing the accuracy of the posteriors obtained using R-VGA-Whittle and HMC-Whittle. All experiments were carried out on a high-end server, with an NVIDIA Tesla A100 80GB graphics processing unit (GPU) used to parallelise over the $S=1000$ Monte Carlo samples needed for estimating the expectations of the gradient and Hessian of the Whittle log-likelihood at each iteration of Algorithm 2.
Sections 3.1 to 3.4 show results for simulated data examples involving a univariate linear Gaussian SSM, a univariate stochastic volatility (SV) model, a bivariate SV model, and an SSM with Student’s $t$ error, where the latent states follow an ARFIMA process. Section 3.5 applies R-VGA-Whittle to fit univariate and bivariate SV models to real data.
@@ 3.1 Univariate linear Gaussian SSM
Table 1: Computing time (in seconds) for R-VGA-Whittle, HMC-Whittle and HMC-exact for each model we consider in this study. The number of observations $T$ for each example is also shown.
$T$ R-VGA-Whittle HMC-Whittle HMC-exact MCMC-stochvol
Linear Gaussian 10000 8 10 348 -
SV (sim. data) 2000 8 4 383 11
2D SV (sim. data) 5000 53 1464 35421 -
ARFIMA (sim. data) 50000 823 14220 - -
SV (real data) 2649 8 3 1508 14
2D SV (real data) 2649 49 1136 37227 -
A comparison of the computing time between R-VGA-Whittle, HMC-Whittle and HMC-exact for this model can be found in Table 1. The table shows that R-VGA-Whittle is slightly faster than HMC-Whittle and more than 40 times faster than HMC-exact.
@@ 3.2 Univariate stochastic volatility model
[…]
KimSV models are a popular class of non-Gaussian financial time series models used for modelling the volatility of stock returns. Under a log-squared transformation of the data, a univariate SV model can be written in a linear additive form, where the error follows a log $\chi^{2}_{1}$ distribution (Ruiz,, 1994). A popular approach for estimating this type of model was proposed by Kim et al., (1998), in which the distribution of the error terms in the log-transformed data is approximated using a Gaussian mixture. This approach can be implemented via the R package stochvol (Kastner and Frühwirth-Schnatter,, 2014); we use this package for comparing to our proposed approach, and refer to this method as MCMC-stochvol.
In this section, we follow Fearnhead, (2011) and consider the following univariate SV model:
@@
We apply a log-squared transformation Ruiz, (1994),
@@
Since $\mathbb{E}(\log(y_{t}^{2}))=\mu$, the parameter $\mu$ can be estimated using the sample mean (Ruiz,, 1994):
@@
$\hat{\kappa}=\sqrt{\exp(\hat{\mu}-\mathbb{E}(\log(\epsilon_{t}^{2})))},$ (30)
[…]
@@
which leads to 95% probability intervals of $(0.57,0.99)$ for $\phi$, and $(0.11,0.44)$ for $\sigma_{\eta}$. The interval for $\phi$ aligns closely with those used in similar studies of the stochastic volatility (SV) model. For example, Kim et al., (1998) employed the transformation $\phi=2\phi^{*}-1$ with $\phi^{*}\sim\textrm{Beta}(20,1.5)$, resulting in a 95% probability interval for $\phi$ of $(0.59,0.99)$. Additionally, Kim et al., (1998) utilised an inverse gamma prior for $\sigma_{\eta}^{2}$ with shape and rate parameters of $2.5$ and $0.025$, respectively, corresponding to a 95% probability interval $\sigma_{\eta}$ for $(0.06,0.24)$, which is comparable to the interval implied by our chosen prior.
The stochvol package uses a Beta$(a_{0},b_{0})$ prior for $\phi$ and a Gamma$(1/2,\lambda_{0}/2)$ prior for $\sigma_{\eta}^{2}$, where $\lambda_{0}$ is the rate parameter. To make the priors comparable between stochvol and R-VGA-Whittle, we tune the hyperparameters $a_{0},b_{0}$, and $\lambda_{0}$ by minimising the sum of squared differences between the 2.5th, 50th and 97.5th quantiles of the prior distributions used in stochvol and those of the prior distribution in R-VGA-Whittle. The resulting prior hyperparameters used with stochvol are $a_{0}=10.777,b_{0}=0.459$, and $\lambda_{0}=19.445$.
We first generate $T=2000$ observations from the univariate SV model given in (24)–(27) with the true parameter values: $\phi=0.99$, $\sigma_{\eta}=0.4$, and $\kappa=2$. We compare the performance of R-VGA-Whittle to the HMC-exact, HMC-Whittle, and MCMC-stochvol methods.
@@
Figure 5 shows the trajectories of the R-VGA-Whittle variational means for all parameters. The 3dB cutoff point for this example is the 10th frequency, after which there are 11 blocks, for a total of 21 updates. Similar to the case of the LGSS model in Section 3.1, the variational mean of $\phi$ is close to the true value from the early stages of the algorithm, while that of $\sigma_{\eta}$ requires more iterations to achieve convergence.
@@
Figure 5: Trajectories of the variational means for the parameters in the univariate SV model fitted to simulated data with $T=2000$, where $\phi=0.99$ and $\sigma_{\eta}=0.4$. The true parameter values are indicated by dashed lines, and the vertical line marks $\tilde{n}$, the point beyond which updates are performed in blocks.
Figures 9(a)9(b) in Section S5 of the online supplement show the HMC-exact, HMC-Whittle, and MCMC-stochvol trace plots for all three methods. In this example, all MCMC methods yield well-mixed chains. However, both HMC-exact and MCMC-stochvol yield posterior samples with higher inefficiency factors, measured as the total number of MCMC iterations after burn-in divided by the effective sample size (ESS), compared to HMC-Whittle, particularly for $\sigma_{\eta}$. This difference arises because HMC-Whittle targets the marginal posterior distribution of the parameters only, whereas HMC-exact samples from the full joint posterior of both the latent states and parameters, which is high-dimensional and more challenging to explore efficiently. Similarly, MCMC-stochvol employs a Gibbs sampling scheme that alternates between sampling the latent states $\{x_{t}\}_{t=1}^{T}$ given the parameters $\bm{\theta}$, and sampling $\bm{\theta}$ given the latent states, which also introduces dependence between successive draws.
@@
Figure 6: Posterior distributions from HMC-exact (blue), MCMC-stochvol (black), and approximate posterior distributions from R-VGA-Whittle (red) and HMC-Whittle (yellow) for data simulated from the univariate SV model with $T=2000$, $\phi=0.99$, and $\sigma_{\eta}=0.4$. Diagonal panels: Marginal posterior distributions with true parameters denoted using dotted lines. Off-diagonal panels: Bivariate posterior distributions with true parameters denoted using the symbol $\times$.
[…]
stochvolWe compare the posterior densities from R-VGA-Whittle, HMC-Whittle, HMC-exact, and MCMC-stochvol in Figure 6. For the parameter $\phi$, the posterior distributions obtained from HMC-Whittle closely resemble those from HMC-exact and MCMC-stochvol, whereas that from R-VGA-Whittle is slightly narrower. The posterior distributions for the parameter $\sigma_{\eta}$ are similar across all four methods.
Table 1 shows that R-VGA-Whittle is a few seconds slower than HMC-Whittle in this example, but faster than MCMC-stochvol, and nearly 50 times faster than HMC-exact.
We now examine the empirical coverage of the 95% credible intervals of R-VGA-Whittle and compare them to those obtained from HMC-exact and HMC-Whittle. We consider different values of $\phi$ in the set $\{0.7,0.8,0.9,0.99\}$, with $\sigma_{\eta}$ now fixed at 0.2. For each setting, we simulate 100 time series, each of length $T=2000$. For each value of $\phi$, the $\hat{R}$ statistic and effective sample sizes averaged across 100 datasets are reported in Table S3 in the online supplement.
Table 2 gives the empirical coverage of the 95% credible intervals for $\phi$ and $\sigma_{\eta}$, over 100 simulated datasets, for different values of $\phi$ and $\sigma_{\eta}=0.2$. The empirical coverages for $\phi$ are generally close to the nominal 95% level across all methods, indicating good calibration. R-VGA-Whittle performs comparably to the HMC-Whittle and HMC-exact methods, although it slightly undercovers $\phi$ when $\phi=0.9$, while HMC-exact and HMC-Whittle maintain more stable coverage across all values of $\phi$. The empirical coverage for $\sigma_{\eta}$ is consistently above 95% across all three methods when $\phi=0.7$, $0.8$, and $0.9$. At $\phi=0.99$, both R-VGA-Whittle and HMC-Whittle exhibit undercoverage, while HMC-exact shows overcoverage. Figure 7 show the comparisons of the 95% credible intervals from R-VGA-Whittle and HMC-exact (left) and R-VGA-Whittle and HMC-Whittle (right) in the case of $\phi=0.9$ for the first 25 simulated datasets. Figures S14S16 in Section S6 of the online supplement show similar plots for $\phi\in\{0.7,0.8,0.99\}$ and $\sigma_{\eta}=0.2$. The figures show that the widths of the 95% credible intervals obtained from R-VGA-Whittle and HMC-Whittle are similar, but slightly different from those obtained with HMC-exact.
(a) HMC-exact and R-VGA-Whittle.
(b) HMC-Whittle and R-VGA-Whittle.
Figure 7: Comparison of the 95% credible intervals from R-VGA-Whittle and HMC-exact (left) and R-VGA-Whittle and HMC-Whittle (right) for the first 25 simulated datasets from the univariate SV model with $T=2000$, $\phi=0.9$ and $\sigma_{\eta}=0.2$. Posterior means are marked with points.
Table 2: Empirical coverage of the 95% credible intervals for $\phi$ and $\sigma_{\eta}$, computed over 100 simulated datasets of length $T=2000$, with $\phi\in\{0.7,0.8,0.9,0.99\}$ and $\sigma_{\eta}=0.2$. Results are reported for R-VGA-Whittle, HMC-Whittle, and HMC-exact.
$\phi$ R-VGA-Whittle HMC-Whittle HMC
Coverage of $\phi$ Coverage of $\sigma_{\eta}$ Coverage of $\phi$ Coverage of $\sigma_{\eta}$ Coverage of $\phi$ Coverage of $\sigma_{\eta}$
0.7 0.96 0.99 0.97 0.99 0.99 0.99
0.8 0.97 0.99 0.99 0.99 0.98 0.99
0.9 0.91 0.99 0.97 0.98 0.96 0.97
0.99 0.94 0.92 0.97 0.89 0.96 0.97
@@ 3.3 Bivariate SV model
Figure 8: Trajectories of the variational means for the parameters in the bivariate SV model fitted to simulated data with $T=5000$. The true parameter values are indicated by dashed lines, and the vertical line marks $\tilde{n}$, the point beyond which updates are performed in blocks.
@@
Figures 10(a)10(c) in Section S5 of the online supplement show the HMC-exact and HMC-Whittle trace plots for this model. HMC-exact samples are highly correlated, especially those of $\Sigma_{\eta_{22}}$. We find that it is necessary to thin the HMC-exact samples by a factor of 50 for the samples to be considered approximately uncorrelated draws from the posterior distribution. In contrast, HMC-Whittle produces well-mixed posterior samples. Section S5 of the online supplement also shows that the inefficiency factors from HMC-Whittle are much lower than those from HMC-exact for all parameters.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle, and HMC-exact in Figure 9. In this example, we observe broad agreement between the three methods, though the posterior densities from R-VGA-Whittle and HMC-Whittle are wider than those from HMC-exact. The posterior density of $\Phi_{22}$ from R-VGA-Whittle puts more probability on values close to one than those from the other two methods, while the posterior density of $\Sigma_{\eta_{21}}$ from both R-VGA-Whittle and HMC-Whittle appear slightly shifted towards zero compared to that from HMC-exact. As shown in Table 1, R-VGA-Whittle is more than 25 times faster than HMC-Whittle, and over 600 times faster than HMC-exact in this example.
@@
3.4 ARFIMA model with t-distributed measurement noise
Autoregressive Fractionally Integrated Moving Average (ARFIMA) models (Granger and Joyeux,, 1980) are a class of long-memory time series models that extend the well-known class of Autoregressive Integrated Moving Average (ARIMA) models (Box et al.,, 2015) to non-integer values of the differencing parameter $d$. Of particular interest is the case $-0.5<d<0.5$, where the ARFIMA process is stationary and has long-range dependence with an autocovariance function that decays very slowly. ARFIMA processes are particularly useful for modeling macroeconomic and financial time series; see, for example,  Bhardwaj and Swanson, (2006) for a comparison of the predictive accuracy (evaluated based on real datasets of stock index returns) of ARFIMA models to other conventional time series models such as autoregressive (AR), moving average (MA), and autoregressive moving average (ARMA) models.
We use the fracdiff R package (Maechler et al.,, 2024) to simulate a time series of length $T=50000$ from a model in which an underlying ARFIMA$(1,d,1)$ process is contaminated by measurement noise following a Student’s $t$-distribution with $\nu$ degrees of freedom:
$y_{t}$ $=x_{t}+\epsilon_{t},\quad\epsilon_{t}\sim t_{\nu}(0,1),$
$\mathbf{\Psi}(L)(1-L)^{d}x_{t}$ $=\mathbf{\Theta}(L)\eta_{t},\quad\eta_{t}\sim N(0,\sigma_{\eta}^{2}),\quad t=1,\dots,T,$ (41)
where $L$ is the lag operator such that $Lx_{t}=x_{t-1}$, $\mathbf{\Psi}(L)=1-\phi L$, $\mathbf{\Theta}(L)=1+\theta L$, and the fractional differencing operator $(1-L)^{d}$ is defined through a binomial series expansion (Hosking,, 1981). We set the true parameters to $\phi=0.3,\theta=0.7,d=0.25,\sigma_{\eta}=1$, and $\nu=4$. This model is computationally challenging to fit using standard likelihood-based methods (Martin and Wilkins,, 1999); estimating parameters in ARFIMA-type models is more computationally feasible in the frequency domain using the Whittle likelihood (Salomone et al.,, 2020). As a result, we only compare R-VGA-Whittle and HMC-Whittle for this example.
The spectral density for $\{y_{t}\}_{t=1}^{T}$ is given by
$f_{y}(\omega_{k})=f_{x}(\omega_{k})+f_{\epsilon}(\omega_{k}),$ (42)
where the spectral density for the underlying ARFIMA$(1,d,1)$ process is given by
$f_{x}(\omega_{k})=\sigma^{2}_{\eta}\left\lvert 1-e^{-i\omega_{k}}\right\rvert^{-2d}\left\lvert\frac{\mathbf{\Theta}(e^{-i\omega_{k}})}{\mathbf{\Psi}(e^{-i\omega_{k}})}\right\rvert^{2},\quad k=-\left\lceil\frac{T}{2}\right\rceil+1,\dots,\left\lfloor\frac{T}{2}\right\rfloor,$ (43)
and the spectral density for the Student’s $t$ noise process with degree of freedom $\nu$ is
$f_{\epsilon}(\omega_{k})=\frac{\nu}{\nu-2}.$ (44)
As in previous examples, we use the following transformations to make the parameters unconstrained: ${\tilde{\phi}=\tanh^{-1}(\phi)}$, $\tilde{\theta}=\tanh^{-1}(\theta),\tilde{d}=\tanh^{-1}(2d),\tilde{\sigma}_{\eta}=\log(\sigma_{\eta}^{2})$, and $\tilde{\nu}=\log(\nu-2)$.
To improve the convergence of HMC-Whittle and R-VGA-Whittle, we construct prior distributions for the parameters based on their maximum-Whittle-likelihood estimates. The prior distribution we use is
$p(\bm{\theta})=q_{0}(\bm{\theta})=\mathrm{Gau}\left(\begin{bmatrix}\tanh^{-1}(\hat{\phi})\\ \tanh^{-1}(\hat{\theta})\\ \tanh^{-1}(2\hat{d})\\ \log(\hat{\sigma}_{\eta})\\ \log(\hat{\nu}-2)\end{bmatrix},\begin{bmatrix}0.25&0&0&0&0\\ 0&0.25&0&0&0\\ 0&0&0.25&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\\ \end{bmatrix}\right),$ (45)
where $\hat{\phi},\hat{\theta},\hat{d},\hat{\sigma}_{\eta}$, and $\hat{\nu}$ are the maximum-Whittle-likelihood estimates of $\phi,\theta,d,\sigma_{\eta}$, and $\nu$. This prior distribution results in $95\%$ probability intervals of $(-0.56,0.86)$ for $\phi$, $(-0.23,0.94)$ for $\theta$, $(-0.21,0.45)$ for $d$, $(0.39,2.76)$ for $\sigma_{\eta}$, and $(2.29,16.98)$ for $\nu$.
For this model, we find it necessary to increase the number of damping observations in R-VGA-Whittle to ${n_{damp}=100}$. All other tuning parameters for R-VGA-Whittle are kept the same as in other examples. We thin the HMC-Whittle posterior samples by a factor of 10 to reduce autocorrelation in the Markov chain. A comparison of the posterior distributions from R-VGA-Whittle and HMC-Whittle is shown in Figure 10. We find that the posterior distributions from both R-VGA-Whittle and HMC-Whittle contain the true parameters, but that the posterior distributions from R-VGA-Whittle are slightly narrower than those from HMC-Whittle.
R-VGA-Whittle takes less than 15 minutes to fit this model, whereas HMC-Whittle takes nearly 4 hours (see Table 1), making R-VGA-Whittle more than 17 times faster in this case. This further highlights the computational efficiency of R-VGA-Whittle for larger datasets. For ARFIMA models, larger datasets are not uncommon. Salomone et al., (2020), for instance, uses an SV model with an underlying ARFIMA process to analyse a time series of one-minute Bitcoin returns with an order of magnitude more observations than the time series in this simulation example.
Figure 10: Approximate posterior distributions from R-VGA-Whittle (red) and HMC-Whittle (yellow) for data simulated from the SSM with Student’s $t$ error, where the latent states follow an ARFIMA$(1,d,1)$ process, with $T=50000$ observations. Diagonal panels: Marginal posterior distributions with true parameters denoted using dotted lines. Off-diagonal panels: Bivariate posterior distributions with true parameters denoted using the symbol $\times$.
3.5 Univariate and bivariate SV models with real data